OpenStudy (anonymous):

Show that the sequence defined by \(a_1=1 \text{ } a_{n+1}=3-\frac{1}{a_n}\) is increasing and \(a_n<3\) for all n. Deduce that \(\left\{ a_n \right\}\) is convergent and find its limit.

OpenStudy (anonymous):

Sorry guys, I have no idea how to do this one... Note, I could have typed it a little better: it is defined as: \[a_1=1 \text{ , } a_{n+1}=3-\frac{1}{a_n}\]

OpenStudy (anonymous):

assuming it has a limit, the limit is easy to find set \[x=3-\frac{1}{x}\] and solve for \(x\)

OpenStudy (anonymous):

it is more or less clear that if it is increasing, then \(3-\frac{1}{a_n}<3\) as you are subtracting a positive number for 3 that means you have an increasing sequence, bounded above by 3, so it must have a limit the hard part is showing that it is increasing

OpenStudy (anonymous):

*from 3

OpenStudy (anonymous):

Well the number that gets subtracted is getting smaller each time, so the overall value gets bigger each time

OpenStudy (anonymous):

ok we can do that one too

OpenStudy (anonymous):

you want to show that the sequence is increasing, meaning that \(a_{n+1}>a_n\) right?

OpenStudy (anonymous):

lets right down specifically what that means since \(a_{n+1}=3-\frac{1}{a_n}\) this means you want to show \[3-\frac{1}{a_n}>a_n\]

OpenStudy (anonymous):

*write down

OpenStudy (anonymous):

@Loser66 for some reason I can't reply to your message. Here it is: Ahh okay that makes sense. I think we should do it like in Cal2 @satellite73 Yes...

OpenStudy (anonymous):

Increasing means that the following term is bigger than the previous one, for all terms. Since a1=1 and a2 will be = 3-1=2, that means that a1<a2. But that doesn't prove the "for all" part, right? Or is that enough?

OpenStudy (anonymous):

no that doesn't prove it for all

OpenStudy (anonymous):

I don't know how... :|

OpenStudy (anonymous):

maybe induction?

OpenStudy (anonymous):

assume that \(a_n<3\) and prove that \(a_{n+1}<3\) as well

OpenStudy (anonymous):

I will read up on that tomorrow morning, can't remember exactly how that works

OpenStudy (anonymous):

show it is the case for \(n=2\) which you just did the assume it is true for \(n=k\) and then show it is true for \(n=k+1\) in other words if you assume \(a_k<3\) prove that \(a_{k+1}<3\) this should not be hard

OpenStudy (anonymous):

since \[a_{k+1}=3-\frac{1}{a_k}\] and we assume that \(a_k<3\) making \(\frac{1}{a_k}>\frac{1}{3}\) this tells us that \(3-\frac{1}{a_k}<3\)

OpenStudy (anonymous):

Thanks! Sorry, I need to get to bed now, otherwise tomorrow will be hell. :) Cya